Ask your own question, for FREE!
Mathematics 11 Online
OpenStudy (anonymous):

A game has three possible outcomes: win, lose, or tie. The probability of a win is 0.4, the probability of a loss is 0.1, and the probability of a tie is 0.5. On a single play of the game the probability of NOT winning is

OpenStudy (anonymous):

maybe around 0.4

OpenStudy (anonymous):

can I have a fan and medal or the question?

OpenStudy (anonymous):

not winning means either tie or lose. So 0.5 + 0.1 = 0.6

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
2 days ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
2 days ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!